OpenStudy (anonymous):

The monthly incomes of A, B and C are in the ratio 3 : 4 : 5 while their savings are in the ratio 1 : 2 : 3. If the savings of A is Rs. 600 per month equivalent to 10% of his income, how much does C spend per month?

OpenStudy (anonymous):

so a income=6000 10/100x=600

OpenStudy (anonymous):

or else 90x/100=600

OpenStudy (anonymous):

so x=2000 then c should be 10000

OpenStudy (anonymous):

a spends 10% of saving which has ratio 1 so for 30 % saving

OpenStudy (anonymous):

so spending of c is 7000?

ganeshie8 (ganeshie8):

lets work A first : savings = 600 = 10% of his income so, income = 6000, right ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

next, their savings are in the ratio 1 : 2 : 3 since A has a saving of 600, that means C must have a saving of 3*600 = 1800, right ?

OpenStudy (anonymous):

yes i have that problem

ganeshie8 (ganeshie8):

next, whats the income of C ?

OpenStudy (anonymous):

so 8200

OpenStudy (anonymous):

8200 he spends,10000 income

ganeshie8 (ganeshie8):

incomes of A, B and C are in the ratio 3 : 4 : 5 since A has an income of 6000, that means C must have an income of 6000*5/3 = 10.000

ganeshie8 (ganeshie8):

so his spendings = 10,000 - 1800 = 8200 you're right :)